### v4

# Conductivity of concentrated salt solutions

Abstract
The conductivity of concentrated salt solutions has posed a real puzzle for theories of electrolytes. Despite a quantitative understanding of dilute solutions, an analytical theory for concentrated ones remains a challenge almost a century, although a number of parameters and effects incorporated into theories increases with time. Here we show that the conductivity of univalent salt solutions can be perfectly interpreted using a simple model that relies on a modified mean-field description of electrostatic effects and on a classical approach to calculating colloid electrophoresis. We derive a compact equation, which predicts that the ratio of conductivity at a finite concentration to that at an infinite dilution is the same for all salt, if it is plotted against a product of the harmonic mean of ion hydrodynamic radii and the square root of concentration. Our equation fits very well the data for inorganic salts (up to a few mol/l), although at a very high dilution the relaxation correction seems necessary.
preprint: APS/123-QED
I Introduction
The electrolyte solutions are ionic conductors, thanks to cations and anions formed as a result of dissociation. The electrical conductivity $K$ is one of their most important trait, which is widely used in chemical, biological as well as other applications, and the performance of many (e.g. energy storage) devices depends entirely on the ionic transport [1, 2, 3]. The amount of water in Earth’s mantle is inferred from the conductivity data [4]. Besides, historically, the conductivity is the most important source of information on electrolyte properties (e.g. ion pairing) [5].
The physical origin of ionic conductivity is more or less clear. If an electric field $E$ is applied, ions migrate relative to a solvent by generating an electric current of density $J=KE$ (Ohm’s law). The migration speeds of ions are given by
$V_{\pm}=M_{\pm}E,$ (1)
where $M_{\pm}$ are their mobilities. The current densities induced in an univalent electrolyte are then $J_{\pm}=en_{\infty}V_{\pm}$, where $e$ is an elementary positive charge and $n_{\infty}=n_{\pm}$ is the number density (concentration) of an electrolyte solution. Consequently, the current density of a solution reads $J=J_{+}-J_{-}=eEn_{\infty}M,$ where $M=M_{+}-M_{-}$ is the difference in mobilities of cations and anions, and
$K=en_{\infty}M$ (2)
Thus, the calculations of the conductivity at a given concentration are reduced to those of $M$.
The quantitative understanding of ion mobilities is a challenging problem that has been addressed over nearly a century and by many groups, which is often termed a central issue of chemical physics [6, 7]. The simplest expression can be derived by postulating the Stokes resistance to the ion propulsion
$M_{\pm}\simeq M_{\pm}^{0}\simeq\pm\dfrac{e}{6\pi\eta R_{\pm}},$ (3)
where $\eta$ is the dynamic viscosity of the solvent and $R_{\pm}$ are the hydrodynamic radii of the cation and anion. $M_{\pm}^{0}$ given by Eq.(3) and termed the mobilities at an infinite dilution do not depend on salt concentration. In this model
$M\simeq M_{0}\simeq\dfrac{e}{3\pi\eta R_{h}}\,\mathrm{and}\,K\simeq K_{0}\simeq\dfrac{e^{2}n_{\infty}}{3\pi\eta R_{h}},$ (4)
where $\eta$ is the dynamic viscosity of a solvent and $R_{h}=2(R_{+}R_{-})/(R_{+}+R_{-})$ is the harmonic mean of hydrodynamic radii. However, experiments on a conductivity show that $M$ and $K$ are generally smaller than predicted by (4) and this discrepancy augments on increasing salt concentration [8, 1]. Chemists have long used this fact to infer a degree of dissociation (or an ion pairing). Physicists might view this simply as unreliability of hydrodynamic arguments based on the Stokes force.
There is a large literature describing attempts to provide a satisfactory theory of electrolyte conductivity. One of the first systematic treatments of the influence of $n_{\infty}$ on conductivity was contained in a remarkable paper by Onsager [9] who clarified a mechanism for the conductivity reduction, which involved ionic electrophoresis and relaxation, and showed that at a very high dilution the correction to $M_{0}$ depends on the square root of the concentration. In efforts to better understand the connection between the ion mobilities and salt concentration many authors extended the theory, but failed to come to grips with the conductivity at high salt [10]. However, most electrolyte solutions in nature and various applications are concentrated. Thus, salt concentration in human blood plasma is about 0.15 mol/l, in the Atlantic Ocean it is $\simeq 0.6$ mol/l, Li-ion batteries and supercapacitor electrolytes are usually of concentration $1-2$ mol/l, reference electrolytes of pH-meters and glass micropipette electrodes - of concentration 3 mol/l.
Recent attempts at improvements on the theory has been focused on a more sophisticated description of electrostatic and hydrodynamic effects. Various techniques, such as the mean spherical approximation [11, 6], mode-coupling theory [12, 13], density functional approach [12, 14], and more [15, 16, 17, 7, 18] have been employed. These publications involve additional parameters and contributions, and mostly rely on numerical calculations. This makes them difficult to use and limits prediction capabilities, but simple analytical formulas applied for concentrated solutions have not yet been obtained, although there have been some attempts to deduce them from microscopic theories. Useful analytical solution is known [19], but expressed in terms of special functions and can hardly be called simple. Avni et al. [14] derived an approximate conductivity expression, but do not recognize its inapplicability at molar concentrations due to non physical behavior of electrophoretic and relaxation terms [20]. Recently we suggested a completely different line to attack the ionic electrophoresis problem [21]. Assuming cations and anions of equal size, we have carried calculations of an electrophoretic term and derived an equation for a corresponding reduction in relative conductivity, $K/K_{0}$, that applies even at a very low dilution. However, this paper has not given sufficient attention to the relaxation issue, which is traditionally invoked in the conductivity theories. It is also, in practice, unlikely that both cations and anions have exactly the same size, so the status of our theory for real salt solutions remains somewhat obscure.
The present paper extends and generalizes the prior analysis [21] to solutions composed of inorganic ions of unequal radius $R_{+}\neq R_{-}$. We derive a simple equation for an electrophoretic contribution to conductivity valid up to a few mol/l and argue that relative conductivities, $K/K_{0}$, of all inorganic salts plotted against $R_{h}\sqrt{n_{\infty}}$ would collapse into a single curve. This conclusion is supported by providing a comparison with data for several standard salts, which shows that our theory is in excellent agreement with experiment up to concentrations of a few molars. We also demonstrate that although the relaxation effect affects a small decrement to a relative conductivity at high dilution, it becomes negligibly small in concentrated solutions, where the reduction in $K/K_{0}$ is significant.
II General considerations
We consider a bulk 1:1 salt solution assuming that the description of its global static properties can be restricted to a mean-field Poisson-Boltzmann theory. The dimensionless electrostatic potential around each ion $\phi(z)=e\Phi(z)/(k_{B}T)$, where $k_{B}$ is the Boltzmann constant and $T$ is a temperature, represents a continuous function that depends on all other ions. The Debye screening length of a solution, $\lambda_{D}=\left(8\pi\ell_{B}n_{\infty}\right)^{-1/2}$, is defined as usually with the Bjerrum length, $\ell_{B}=\dfrac{e^{2}}{\varepsilon k_{B}T}$, where $\varepsilon$ is the solvent permittivity. By analyzing the experimental data it is more convenient to use the concentration $c_{\infty}[\rm{mol/l}]$, which is related to $n_{\infty}[\rm{m^{-3}}]$ as $n_{\infty}\simeq N_{A}\times 10^{3}\times c_{\infty}$, where $N_{A}$ is Avogadro’s number. The Bjerrum length of water at $T\simeq 298$ K is equal to about $0.7$ nm leading to a useful formula for 1:1 electrolyte
$\lambda_{D}[\rm{nm}]\simeq\frac{0.305[\rm{nm}]}{\sqrt{c_{\infty}[\rm{mol/l}]}}$ (5)
Thus upon increasing $c_{\infty}$ from $10^{-5}$ to $3$ mol/l the screening length is reduced from about 100 down to 0.18 nm.
Figure 1: Sketch of the cation and anion with hydrodynamic radii $R_{+}$ and $R_{-}$ in the bulk electrolyte solution characterized by $\lambda_{D}$. The cation propels with the speed $V_{+}$ in the direction of the electric field $E$. Its zeta potential $\zeta_{+}$ (or dimensionless mobility) is positive. The anion of negative $\zeta_{-}$ migrates against the field with the velocity $V_{-}$.
Table 1: Some typical values of hydrodynamic radii of ions [22], they harmonic means and ratios.
Salt $R_{+}$, nm $R_{-}$, nm $R_{h}$, nm N
KBr 0.1295 0.1179 0.123 1.1
NaCl 0.184 0.1245 0.148 1.48
LiI 0.238 0.1135 0.154 2.1
We recall that the mean-field theory neglects correlations and finite sizes of ions. By contrast, to describe the dynamic response to an external field, the ion hydrodynamic radii $R_{\pm}$ should be taken into account. Inorganic ions have hydrodynamic radii from 0.1 to 0.3 nm [22], and we present their values for some univalent electrolytes in Table 1, together with harmonic means $R_{h}=2R_{+}R_{-}/(R_{+}+R_{-})$ and $N=R_{+}/R_{-}$. These apparent contradictions indicate that in reality, ions are neither point-like nor particles of finite size, exhibiting the properties of the former or the latter only depending on the circumstances. This can be referred to as the point-particle duality of ions (by analogy to the familiar concept of wave-particle duality of electrons in quantum mechanics). Such a duality can be introduced, if an electrolyte solution is modelled as a colloidal suspension of so-called central ions representing isolating particles of radius $R_{\pm}$ with a constant surface charge density ±$\pm e/(4\pi R_{\pm}^{2})$ [21]. Such a sphere induces the same outer field as a point charge ±$\pm e$ located at its center would, and the salt-dependent surface potential $\phi_{s\pm}$ is established self-consistently. By contrast, the electrostatic diffuse layers (EDLs or clouds) are treated on the classical mean-field level and formed by point-like ionic species, thus their extension is defined by $\lambda_{D}$. Two further comments should be made: (i) Since the model prevents point-like EDL ions from penetrating inside the sphere, the need to invoke an additional electrostatic cut-off [14] is removed. (ii) The approach neglects electrostatic correlations inside the cloud, but provides positional correlations between central ions, although on average they are randomly distributed.
Since the electrostatics of ions involves two length scales, it is convenient to introduce their dimensionless radii $\varrho_{\pm}=R_{\pm}/\lambda_{D}$. Note that in the range of $c_{\infty}$ below 3 mol/l we consider here, the values of ϱ±subscriptitalic-$\varrho_{\pm}$ remain smaller than unity or very close to it. Say, if we set $c_{\infty}=2$ mol/l, then for a largest ion (Li+) in Table 1 we obtain $\varrho_{+}\simeq 1$, but for a smallest (I-) we get $\varrho_{-}\simeq 0.5$.
When an electric field $E$ is applied, an electro-osmotic flow around central ions, referred below simply to as ions, is induced. The electroosmosis takes its origin in the EDL, where a tangential electric field generates a force that sets the fluid in motion. The emergence of this flow in turn provides hydrodynamic stresses that cause the propulsion of ions with a velocity given by Eq.(1) as sketched in Fig. 1. The space between clouds of two ions is electro-neutral, and this region is force-free (no flow) provided they do not overlap. Thus, the problem of ionic electrophoresis can be treated on a single-particle level.
We return to the importance of the relaxation later, by focusing now on electrophoresis solely. This contribution to a conductivity is traditionally calculated assuming that the surrounding cloud remains undisturbed during the migration. The electrophoretic mobilities of these ions can be defined using the known expression for an electro-osmotic mobility [23, 24] taken with the opposite sign
$M_{\pm}=\dfrac{\varepsilon Z_{\pm}}{4\pi\eta}=\dfrac{e}{4\pi\eta\ell_{B}}\zeta_{\pm},$ (6)
where $Z_{\pm}$ are the electrokinetic or zeta potentials of cations and anions, and $\zeta_{\pm}=eZ_{\pm}/k_{B}T$. In such a definition the geometry factor is hidden inside $\zeta$. Note that by introducing the dimensional velocity $v_{\pm}=\dfrac{4\pi\eta\ell_{B}}{eE}V_{\pm}$, one can rewrite (1) as $v_{\pm}=\zeta_{\pm},$ which points clearly that the electrophoretic speed of ions is set by their zeta potentials.
III Results and discussion
Zeta potentials of small inorganic ions are given by [21]
$\zeta_{\pm}\simeq\phi_{s\pm}\mathcal{F_{\pm}}\simeq\pm\dfrac{\ell_{B}}{R_{\pm}(1+\varrho_{\pm})}\mathcal{F_{\pm}},$ (7)
where $\mathcal{F_{\pm}}$ are the special functions derived by Henry [25]
±$\mathcal{F_{\pm}}=1-e^{\varrho_{\pm}}\left[5\operatorname{Ei}_{7}(\varrho_{\pm})-2\operatorname{Ei}_{5}(\varrho_{\pm})\right].$ (8)
[…]
±$\mathcal{F_{\pm}}\simeq\dfrac{2}{3}\left[1+\left(\dfrac{\varrho_{\pm}}{4}\right)^{2}\right]$ (9)
[…]
$\zeta_{\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}(1+\varrho_{\pm})}.$ (10)
Clearly, the upper bound for the zeta potential is attained when $\rho_{\pm}\to 0$ (an infinite dilution)
$\zeta_{\pm}\to\zeta_{0\pm}\simeq\pm\dfrac{2\ell_{B}}{3R_{\pm}}.$ (11)
Figure 2: Zeta potentials $\zeta_{\pm}$ computed for K+, Li+, Cl-, I- (solid curves from top to bottom). Dashed lines are obtained using Eq. (11). Symbols show calculations from Eq. (10).
Figure 2 includes computed zeta potentials of several ions. For these examples we take from Table 1 cations (K+, Li+) and anions (I-, Cl-) of smallest and largest radius. Calculations are made using the first equality in (7) with $\phi_{s\pm}$ found numerically from the nonlinear Poisson-Boltzmann equation and $\mathcal{F_{\pm}}$ calculated from Eq. (8). The straight lines corresponding to $\zeta_{0\pm}$ given by (11) describe perfectly the distinct plateau regions at very low concentrations. On increasing concentration, however, the absolute values of $\zeta_{\pm}$ decrease. Also included are calculations from (10). It can be seen that their fit to numerical data is extremely good in the whole concentration range.
Refer to captionRefer to captionRefer to caption
Figure 3: $K_{\infty}/K_{\infty}^{0}$ as a function of $c_{\infty}$ for KBr, NaCl and LiI aqueous solution calculated from Eqs. (13), (14) and (15) (solid, dash-dotted and dashed curves). Open and filled circles indicate experimental data from [27] and [28], open and filled squares show data from [29] and [30].
From Eq.(6) it follows that
$M\simeq\dfrac{e}{4\pi\eta\ell_{B}}(\zeta^{+}-\zeta^{-})$ (12)
Substituting (10) and performing standard calculations we derive
$\dfrac{M}{M_{0}}\equiv\dfrac{K}{K_{0}}\simeq\dfrac{1}{N+1}\left(\dfrac{1}{1+\varrho_{+}}+\dfrac{N^{2}}{N+\varrho_{+}}\right),$ (13)
where $M_{0}$ and $K_{0}$ are given by Eqs. (4). Figure 3 shows $K/K_{0}$ calculated from Eq. (13) as a function of $c_{\infty}$. The calculations are compared with experimental data for three standard salts [27]. As seen from Table 1, for these examples $N$ varies from 1.1 for KBr, but can be larger than 2 (e.g. for LiI). An overall conclusion from this linear scale plot is that the theoretical curve for $K/K_{0}$ is well consistent with experiment. At large concentrations Eq.(13) predicts either slightly lower (KBr), but still fits data quite well. When $N=1$, we get $\varrho_{h}=R_{h}/\lambda_{D}=2\varrho_{+}/(1+N)=\varrho_{+}$, and Eq. (13) reduces to
$\dfrac{K}{K_{0}}\simeq\dfrac{1}{1+\varrho_{h}}.$ (14)
Eq. (14) predicts a monotonic decrease in relative (electrophoretic) conductivity with $\varrho$ as it should be, and $K/K_{0}\to 0$ as $\varrho\to\infty$. For $\varrho_{h}\ll 1$ it reduces to the famous Onsager formula for the electrophoretic effect
$\dfrac{K}{K_{0}}\simeq 1-\varrho_{h}.$ (15)
The calculation from Eq.(14) is also included in Fig. 3. It can be seen that deviations from Eq. (13) are extremely small. The difference between (13) and (14) is always below a few %. We thus conclude that compact approximate Eq. (14), which is easy to handle, can safely be employed to interpret data for inorganic salts or for a predictive purpose.
It follows from (5) that $\varrho_{h}\simeq 3.379R_{h}\sqrt{c_{\infty}}$ for a fixed $T=298$ K, where $R_{h}$ should be taken in nm. This implies that the values of $K/K_{0}$ for any univalent salt solution plotted against (dimensionless) $R_{h}\sqrt{c_{\infty}}$ should collapse into a single curve. That this is indeed so is demonstrated in Fig. 4, where we plot in a semi-log scale a relative conductivity calculated from (14) vs. $R_{h}\sqrt{c_{\infty}}$ along with the experimental data for a variety of inorganic salts at $T=298$[27, 29, 30, 31, 28]. Besides salts presented in Table 1 we include KCl, LiCl, and LiClO4 using their $R_{h}$ = 0.127, 0.163, and 0.174 nm, correspondingly, to re-scale the concentration [22]. Note that since the theory is justified provided $R_{h}\sqrt{c_{\infty}}$ is below ca. 0.305, we employ for this plot only data for $c_{\infty}\leq 3$ mol/l. One can see that the approximate theory is in a good agreement with the data for all salts, perhaps with only insignificant discrepancy at low dilution that is probably comparable with the experimental error. The discrepancy is always in the direction of smaller $K/K_{0}$ than predicted by Eq. (14), but the data for KBr and KCl show slightly larger $K/K_{0}$. The reason for qualitative differences only for these two salts is unclear and is open for discussion, but it must be remembered that this was a first-order calculation only, and we did not expect it to be very accurate. The curve calculated from Eq. (15) is also included in Fig. 4. Clearly, the data obtained at high salt are irreconcilable with this linear equation.
Figure 4: $K/K_{0}$ as a function of $R_{h}\sqrt{c_{\infty}}$ plotted in a semi-log scale (solid curve) and experimental data for standards salts taken from Refs. [27, 29, 30, 31, 28] shown by filled symbols. Dashed-dotted and dashed curve corresponds to calculations from Eqs. (14) and (15).
Figure 5: A decrement to a relative conductivity, $1-K/K_{0}$, plotted against $R_{h}\sqrt{c_{\infty}}$ in a log-log scale (solid curve). Symbols show the same data as in Fig. 4. Dashed-dotted, dashed and dotted lines are calculated from Eqs. (14),  (15) and (16).
To examine the possible discrepancy from data more closely, the results for salts from Fig. 4 are reproduced in Fig. 5, where a decrement to a relative conductivity, $1-K/K_{0}$, is plotted against $R_{h}\sqrt{c_{\infty}}$ in a log-log scale. It can be seen that in dilute solutions, where the decrease in relative conductivity is so small that was even not discernible on the scale of Fig. 4, the data are slightly above the theory indicating some additional retardation in the ion migration. An explanation can be obtained if we invoke the relaxation effect, which is due to an emergence of an additional (retarding) field caused by a distortion of the EDL symmetry during the migration. Onsager showed that when $\varrho_{h}\ll 1$ the electrophoretic and relaxation terms are summed up [9], which in our notations yields
$1-\dfrac{K}{K_{0}}\simeq\varrho_{h}+\frac{1}{3}\left(1-\frac{1}{\sqrt{2}}\right)\frac{\ell_{B}}{\lambda_{D}}.$ (16)
Indeed, Eq. (16) provides an excellent fit to the data up to $R_{h}\sqrt{c_{\infty}}\simeq 6\times 10^{-3}$ (millimolar concentrations). Upon increasing concentration further the experimental data begin to approach to calculations from Eq. (14), and in concentrated solutions, $R_{h}\sqrt{c_{\infty}}\geq 8\times 10^{-2}$ (or $c_{\infty}\geq 0.5$ mol/l), our theory that includes an electrophoretic retardation solely is in excellent agreement with experiment. It is natural to conclude that when dilution is low the relaxation effect reduces with $c_{\infty}$ and disappears at high concentrations. Experimental observations [32, 33], simulations [6], and (numerical) calculations based on a nonlocal electrostatics approach [18] produce an endorsement of our conclusion. We also recall, that Overbeek [34] argued that, although in highly dilute solutions the relaxation effect grows with $c_{\infty}$ (see Eq. (16)), after reaching a maximum this correction to electrophoresis of colloid particles should reduce and disappears at high salt. Exactly the same was found here for inorganic salt solutions, which supports the concept of point-particle duality of ions.
IV Concluding remarks
[…]
  • The electrophoresis of inorganic ions can be accommodated within a theoretical framework that employs only a harmonic mean of hydrodynamic radii of ions and relies on a modified mean-field description of electrostatics (a concept of point-particle duality) and on a classical approach to colloid electrophoresis. The need to invoke specific constants as (arbitrary) parameters, or to correct an ionic concentration is thereby removed;
  • The conductivity of concentrated salt solutions is dictated solely by an ionic electrophoresis. Although the electrophoretic retardation dominates in highly dilute solutions too, to arrive at a more accurate description of their (in fact, extremely small) conductivity reduction, it seems necessary to include the relaxation correction.
This bears on the whole question of what we mean by a degree of dissociation that is often inferred from the conductivity measurements [36]. Our results show that the conductivity is interpreted without invoking the formation of ionic pairs, thus supporting the notion of complete dissociation of strong electrolytes. However remote from mainstream thinking this conclusion may seem, it would be worthwhile to recall that there are still lingering doubts about the reality of ion pairing, at least for univalent electrolytes in high permittivity solvents [5, 37].
Our considerations can be extended to asymmetric multivalent salts. The same concerns the temperature dependence of $K/K_{0}$, which follows from our theory, but requires the validation in terms of fit to experimental results. Another fruitful direction could be to consider the salt-dependence of a mobility of adsorbed ions [38, 39], which impacts electrokinetics [40].
On a more fundamental level, it would be interesting to revisit such issues as the diffusion of ions, effective permittivity and viscosity of salt solutions that can be quantitatively discussed using the standpoint taken here. If all these can be interpreted in a different from a common viewpoint way, the implications are large. A possible result would probably demand a revision of old dogmas and may represent a step forward.
Acknowledgements.
This research was supported by the Ministry of Science and Higher Education of the Russian Federation. We are indebted to E .S. Asmolov, R. Buchner, Y. A. Budkov, G. T. Hefter, B. Rotenberg, and G. A. Tsirlina for feedback and advises.